Calculate [OH-], pH and pOH of a 0.175M KHCO3 solution at 25 degrees celsius

*WRITE IN PRINT (LEGIBLY)*
Answer #1

KHCO3 can acts as a both acid & base. It is amphoteric in nature.

HCO3- is an intermediate form of an carbonic acid and carbonate ion.

We know that, for intermediate form, [H+ ] = K 1 K 2 F + K 1 K w / K 1 + F

From literature ,dissociation constants of carbonic acids are K 1 = 4.46 10 -07 & K 2 = 4.69 10 -11

[H+ ] = (4.46 10 -07) ( 4.69 10 -11 ) 0.175 + ( 4.46 10 -07 ) ( 1.00 10 -14 ) /  4.46 10 -07 + 0.175

[H+ ] = 3.6605 10 -18 + 4.46 10 -21 / 0.1750

[H+ ] = 2.094 10 -17

[H+ ] = 4.58 10 -09 M

We have relation,  [H+ ] [ OH - ] = 1.00 10 -14

[ OH - ] = 1.00 10 -14 / [H+ ] = 1.00 10 -14 / 4.58 10 -09 = 2.18 10 -06 M

We have relation, p H = -log [H+ ] = - log 4.58 10 -09 = 8.34

We have relation, p OH = -log [ OH - ] = - log 2.18 10 -06 = 5.66

ANSWER : [H+ ] = 4.58 10 -09 M

[ OH - ] = 2.18 10 -06 M

pH = 8.34

pOH = 5.66